Aug 14, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m. Eastern

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Tentative Agenda

  1. Announcements
  2. Action Items
  3. FCAL HDFCAL log book
  4. BCAL HDBCAL log book
  5. Reconstruction and Simulation
  6. Calibration
  7. Any other business


Attendees: Elton, Manuel, Will, Laura, Shayne, Mark (JLab); Tegan, Zisis, Andrei (UofR); Sean (NW).

  1. Announcements
    1. Commissioning run: accelerator will start on Sep 19. commissioning of the electron beam to the tagger dump is scheduled to be completed by the beginning of October. First photon beam commissioning should begin first week of October.
  1. Action Items
  2. FCAL HDFCAL log book
    1. Replacing bases
      • Replaced bases two different times. Placing standard information in the logbook that can be used to track/diagnose problems. This information can be passed on to Paul Smith at IU for diagnosis in fixing the problems.
      • Checked that the latest set of replaced bases are still working correctly.
      • Some problems are causing the base HV to trip unexpectedly and needs to be investigated.
      • There is also instability in the readback of HV and currents from the bases which can cause false alarms
      • Investigating how to alarm on trips and at the same time minimize nuisance alarms. Mark: Most alarm system can check for persistence of signals before alarming.
    2. EPICs GUIs
      • GUIs for base setting and monitoring are being scripted.
      • Working with Hovanes to integrate the monitoring into the GlueX slow controls system.
    3. DAQ is not working
      • Triggering is not functioning correctly. Serguei is investigating the problem. Could be due to similar problems with the BCAL after TI boards were updated.
    4. Auxiliary electronics
      • NIM electronic systems will need to be checked to comply with electrical code requirements for the Hall. Need to work with Fernando to integrate some of the signals into the base equipment.
    5. New postdoc
      • Adesh Subedi has accepted an offer to work for IU and will be stationed at JLab upon completing his thesis.
  3. BCAL HDBCAL log book
    1. Mark: work on TDCs
      • Checking connections, found many issues, most corrected. Some remaining including a ribbon cable that needs swapping out. This will be done this afternoon. Technical staff helped with some of the modules that needed lift access.
      • This afternoon we will scan over LED pulses to calibrate the threshold.
    2. Cosmic-rays
      • More 'tandem' data were taken last night. Will: the signals from the top trigger counters were split and added to the data stream. Mark checked that the signals are appearing in the data.
      • Andrei: Inquired about root files. Mark: Most ROOT files are in one place. Mark will analyze them all systematically with a uniform script.
      • Irina and Andrei have simulated the response to cosmic-ray muons and posted some plots in the e-log
      • There was considerable discussion regarding the plots and what to expect.
  4. Magnet: Solenoid is cold and testing is mostly complete. Ready to ramp up to 800 A this afternoon.
    • Need to plan to test of the effect of magnetic fields on the detectors during this operation.
    • Bottom BCAL trigger counters may be most sensitive.
    • FCAL should also check that especially the inner pmts are unaffected.
    • We need to get a schedule from Zisis. (Note added: Plans are tentative. Yi said that they will ramp the magnet to 800 A this afternoon and then ramp down and turn off for the evening. After consulting with George, they might ramp the magnet back up to 1200 A tomorrow. The magnet will be off over the weekend. Early next week it will be brought back to 1200 A and left there for many days.)
  5. Reconstruction and Simulation (Tegan)
    • Working with Justin Stevens MIT to select photons for pi0 reconstruction using BDTs. Initial analysis shows promise but still optimizing
    • Elton: Estimate number of events needed to calibrate the BCAL using this software for use in planning for data taking this fall.
  6. Calibration
  7. Any other business